OpenStudy (anonymous):

What conic section is the following equation? 7y^2+28=14x^2

OpenStudy (anonymous):

HI :

OpenStudy (anonymous):

hi

OpenStudy (jdoe0001):

\(\large 7y^2+28=14x^2\implies \begin{array}{llll} 7y^2+28&-14x^2=0\\ \uparrow&\uparrow\\ \textit{positive }y^2&\textit{negative }x^2 \end{array}\) what do you think?

OpenStudy (anonymous):

superhyperpola :

OpenStudy (anonymous):

\[14x ^{2}-7y ^{2}=28\] divide by 28 \[\frac{ x ^{2} }{2 }-\frac{ y ^{2} }{ 4 }=1\] which is of the type \[\frac{ x ^{2} }{a ^{2} }-\frac{ y ^{2} }{b ^{2} }=1\] which is a hyperbola.

OpenStudy (anonymous):

i was thinking hyperbala but im just confused on how you tell the diference

OpenStudy (ranga):

Try putting x = 0 in surjithayer's equation above. Will there be a y-intercept?

OpenStudy (anonymous):

\[Here a=\sqrt{2},b=2\]

OpenStudy (ranga):

This hyperbola has x-intercepts and no y-intercept. It is a horizontal hyperbola.

OpenStudy (ranga):



OpenStudy (anonymous):

thanks

OpenStudy (ranga):

yw.